# The Rise of Ionized Gas Filaments in Early-Type Galaxies

### v1

@@ 4 Results
The MUSE data reveals 54/126 () non-central ETGs with H emission in their spectrum, tracing warm ionized gas. These nebulae are divided into two general categories, based on their morphology and kinematics: rotating disks and filamentary networks. The histogram presented in Fig. 1 shows the count and mass distribution of these two morphological classifications. The full stellar mass range of all sources with H emission span 8.25 log(M⋆/M⊙) 11.02. The warm gas morphology and stellar mass of all 54 non-central ETGs are provided in Table. LABEL:tab:results.
@@ 4.1 Rotating warm gas disks
The rotating gas disks dominate the H morphological distribution in our sample. We find that 35 out of 54 H-emitters exhibit rotating disks, which we define as warm nebulae showing a clear velocity gradient indicative of ordered rotation. The sizes of the gas disks, defined as the projected length of the disk’s semi-major axis, range from roughly 0.1 - 8 kpc.
@@ 4.2 Warm gas filaments
The remaining 19 non-central ETGs with H emission have warm gas morphologies in the form of filamentary networks. A source is classified as filamentary if its gas nebulae show compact or extended, irregular structures–typically extending from the optical center of the galaxy. Additionally their gas velocity fields are generally absent of clear, ordered rotation.
The optical continuum, H flux, velocity, and velocity dispersion for all 19 filamentary sources in our sample are shown in Fig. 3. To quantify the size of the filaments, we refer to the projected distance from the optical center of the galaxy to its most radially extended filament. The size of the filaments range from 0.3 - 4 kpc.
@@ 4.3 Stellar kinematics
Using the GIST pipeline (see Sec. 3.1.1) we derive the line-of-sight velocity and velocity dispersion maps tracing the stellar population. of our ETGs with H emission show regular, ordered rotation in their stellar velocity fields. Low-mass (log(M⋆/M⊙) 9.5) sources occupy of the ETGs that are absent of ordered stellar rotation. For some of these low-mass ETGs, this is likely attributed to their low stellar-surface brightness preventing the pipeline from adequately deriving the galaxy’s underlying line-of-sight velocity distribution.
The H and stellar velocity fields of ETGs hosting rotating H disks in general show kinematic and spatial overlap between their gaseous and stellar disks. This is not the case for the H filamentary sources. The asymmetric, non-rotating filamentary nebulae in nearly all sources appear to be dissociated with their stellar population. Sec. 5.1.1 and 5.2.4 discuss the implication that coupling (or decoupling) between the gaseous and stellar components have on interpreting the internal physics driving the host galaxy’s evolution.
@@ 4.4 Hot gas content
In Fig. 4, we compare the diffuse X-ray luminosity in 0.5-2.0 keV (normalized by stellar mass) against the stellar mass for all ETGs in our sample with available MUSE and Chandra data. The horizontal dashed line provides the average X-ray emissivity of unresolved LMXBs from Hou et al. (2021). Galaxies sitting above this line have soft diffuse X-ray emission exceeding what is expected from unresolved LMXBs and likely contain a hot gas reservoir. We distinguish our data by those with filamentary ionized gas (red x’s), ionized gas in the form of rotating disks (blue circles), and those where H went undetected (gray diamonds). The inverted filled triangles show filamentary systems where the emissivity was computing using an upper-limit on . Inverted open triangles show filamentary systems not observed by Chandra, and their upper-limit on from eROSITA observations is gathered from eRASS1 (Predehl et al., 2021; Tubín-Arenas et al., 2024). All ETGs with filamentary ionized gas are likely to contain hot gas, as are a strong percentage of the ETGs containing rotating gas disks. About half of the ETGs where H went undetected in emission contain hot gas. Therefore, we observe a mild connection between the hot gas content and the morphology of the warm gas, when compared at nearly the same mass range.
@@ 4.5 A case study with NGC 4374
We further derived the electron pressure of each of the two phases for the same regions.
To derive the pressure of the warm phase, we adopted the commonly used value of 104 K for the electron temperature () at each pixel, since not all of the [NII] auroral lines required for extracting the warm gas temperature are detected. To compute the warm gas electron number density () at each spaxel within its MUSE image, we utilized PyNeb (Luridiana et al., 2015) to estimate this value based on a given spaxel’s [SII]6717/[SII]6731 value. To determine the electron temperature and number density of the hot gas, we extracted the X-ray spectra from these regions (using the same local background described earlier). We then fit the spectra with an absorbed thermal plasma model, phabsapec, in Xspec version 12.12.1 (Arnaud, 1996). We derive of the hot gas using the fit’s reported normalization parameter, , via:
@@ 5.1.1 Stellar mass loss
Processes involving an ETG’s stellar population can build up its warm ISM, specifically continual mass loss from post-asymptotic giant branch stars (pAGB) and planetary nebulae (Parriott & Bregman, 2008; Bregman & Parriott, 2009; Voit & Donahue, 2011). Material that is gradually shed from stars should be mostly shock-heated and mix in with the ambient, hot atmosphere. However, numerical hydrodynamics suggest that under the right conditions pertaining to the density of the hot ambient medium, as well as a star’s velocity through that medium, can allow the shed material to dodge this shock heating thus preventing it from entering the hot phase (Parriott & Bregman, 2008). In this case, small, but non-negligible amounts of stellar mass loss will remain cool and be distributed within the ISM.
If stellar mass loss is the dominant origin of warm gas, one would expect there to be spatial and kinematic overlap between the stars and gas. This can be observed by comparing the velocity fields of both populations. To do this, we calculate the kinematic position angle (PA; Krajnovic et al. (2006)) of both the stars and gas. The PA allows for a consistent, geometric interpretation of the bulk motion of the stars and gas. Specifically, it is defined as the angle sweeping from the (projected) north to the location of the maximum positive velocity, in the counterclockwise direction.
@@
Using the threshold of from Lagos et al. (2015) to classify sources as having decoupled stars and nebulae, we find that all filamentary sources, except for ESO 380-027, show kinematic misalignment – unlike the majority of the rotating disks. This corroborates our analysis in Eskenasy et al. (2024), where we found that all non-central ETGs with warm filaments had widely offset PAs between the gas and stars. This frequently observed misalignment observed in ETGs with warm filaments suggests that the two populations are decoupled from each other, hinting that stellar mass loss is likely a subdominant process in forming their warm gas.
@@ 5.1.2 Cooling from the hot atmosphere
In a recent study of cool-core filaments, a close correlation has been identified between their X-ray surface brightness and H surface brightness – SBX SB (Olivares et al., 2025). Strikingly, the slope of this relation matches exactly that found for the X-ray and H surface brightnesses of stripped tails of jellyfish galaxies (Sun et al., 2022), demonstrating that the multiphase gas in both settings may share the same origin. The formation mechanism of the warm gas can occur in a broader astrophysical setting, and the presence of X-ray emitting hot gas is indispensable in this process.
As shown in Fig. 4, all filamentary ETGs contain hot gas, while not all rotating disk sources or those that lack an H detection contain hot gas. This may imply that the presence of hot gas is a necessary but not sufficient condition for the formation of filamentary warm gas. We overlay NGC 4477 in Fig. 4 (black square) as it is the smallest known system with X-ray cavities and cool-core signatures (Li et al., 2018). Although it was not observed with MUSE, it serves as a reference point demonstrating that the filamentary ETGs within our sample possess a higher hot gas content than NGC 4477.
@@
from Eq. 1 in Nesvadba et al. (2007) assuming case B recombination Osterbrock & Bochkarev (1989). is the H luminosity (in units of erg s-1) and was extracted from the MUSE H flux maps. is the electron density in units of 100 cm-3. To measure the electron density, we used PyNeb Luridiana et al. (2015) which provides the number density as a function of the [SII]6717 and [SII]6731 flux ratio. Here it is assumed that the gas has an electron temperature of K. Note that variations in the gas temperature are expected, and will certainly affect the derived . However, this is a suitable method for a rough estimate, given that direct temperature measurements from auroral lines are generally unfeasible for our systems, due to their faintness.
@@
adopted from Su et al. (2017). Here, is the average particle weight and is the proton mass. is the bolometric X-ray luminosity, which we estimated using our derived 0.5-2.0 keV luminosities provided in Table. LABEL:tab:results and the energy conversion factor from Xspec version 12.12.1. We assume keV for all sources.
@@
Fig. 7 shows the ionized gas masses of all filamentary sources and the time required to produce such reservoirs via cooling. The crosses and inverted filled triangles show filamentary systems where was derived using Chandra detections and upper limits, respectively. Inverted open triangles show filamentary systems not observed by Chandra, and their upper-limit on is gathered from eRASS1. Our derived timescales span years, all less than the Hubble time. These estimates suggest that it is indeed possible for the filamentary warm gas reservoirs in our non-central ETGs to be produced through cooling.
@@
To further investigate the cooling mechanism in non-central ETGs and its connection to BCGs and jellyfish galaxies, we performed an in depth study of NGC 4374, as it is the closest analogy of cool-core BCGs in our sample (see Sec. 4.5). We compared the X-ray to H surface brightness of the nine regions, finding a median value of , which is larger than the BCG filaments and tails of jellyfish galaxies, differing by 9.06. As described in Sec. 4.5, we derived the pressure of each of the two gas phases and found on that on average, the warm filaments, especially those extended to large radii, have higher electron pressure than that in the hot filaments. This result is opposite to what was found in Olivares et al. (2025) for multiphase filaments in cool-core BCGs. Given that the hot and warm gas should be roughly in pressure equilibrium with each other, Olivares et al. (2025) posited that the warm filaments have additional non-thermal pressure components. In the case of NGC 4374, additional pressure supported is needed in the X-ray filaments.
@@
We investigate the role of non-thermal components such as turbulence and magnetic fields, that may be adding pressure support to NGC 4374’s hot ISM. To estimate the turbulent pressure, we refer to Ogorzalek et al. (2017) which computed the 1D turbulent velocity measurements of NGC 4374 and other massive ellipticals through line broadening and resonant scattering measurements. Interestingly, NGC 4374 was found to have one the lowest 1D velocity dispersions in their sample, with an upper limit of km s-1. The turbulent pressure () can be calculated with where is the gas mass density333 where is the mean molecular weight per electron. is the electron number density estimated in Eq. 1. is the proton mass. and , assuming isotropic turbulence. This would bring the average hot gas turbulent pressure among the regions we consider to be dyn cm-2.
@@
The magnetic pressure () can be estimated as where is the magnetic field amplitude. We refer to Finoguenov & Jones (2001) for the estimate of . In their study, they estimated the strength of the magnetic field by determining the Faraday rotation from radio observations, as well as the electron number density of the hot gas which was extracted from the hot gas’s three dimensional density profile using Chandra X-ray observations. They find that the magnetic field associated with the X-ray plasma is particularly weak at G. This corresponds to dyn cm-2. Our estimates of the non-thermal pressure support added by turbulence and magnetic fields in the hot gas still does not account for its pressure imbalance with the warm gas. We posit that an additional non-thermal pressure component. For example, cosmic rays accelerated by NGC 4374’s prominent AGN jet (Jenkins et al., 1977; Finoguenov & Jones, 2001) may be present. Although, the degree to which cosmic rays add pressure support is not well known.
@@ 5.1.3 Mergers and external accretion
Kinematic misalignment between an ETG’s warm gas and stars can be a key tracer of warm gas having an external origin such as galaxy mergers, tidal stripping from satellites, and accretion. Davis et al. (2011) found that it is common for local ETGs to exhibit this misalignment, thus demonstrating that a non-negligible fraction of warm gas in our sample of non-central ETGs is likely to be expected444Although subtle, but key differences in the merger histories -particularly timescales - arise from ranging environments, stellar masses, and whether an ETG is a fast or slow rotator.. Our comparison of the kinematic position angles in Fig. 6 shows nearly all of our filamentary systems have decoupled stars and gas suggesting an external origin of their warm gas.
@@ 5.2 Powering of line emission
The traditional optical-line diagnostic diagram proposed by Baldwin, Phillips and Terlevich (BPT; Baldwin et al. (1981)) is widely used to distinguish ionization sources in galaxies. However, it is known to have a variety of systematic uncertainties which can lead to ambiguous classification of these ionizing sources. Here, we adopt a modified optical-line diagnostic diagram from Ji & Yan (2020) to determine the ionization mechanisms in our sample. This diagram is a reprojected version of the BPT diagram, which ensures consistency of the model predictions among all three line ratios: [OIII]/H (R), [NII]/H (N), and [SII]/H (S). The axes of this diagram are linear combinations of these three line ratios given as
@@
Fig. 8 shows our systems placed in this reprojected diagram. As described in Ji & Yan (2020), spaxels that fall to the left suffer from less than 10 percent AGN contamination, while those to the right suffer from greater than 10 percent AGN contamination. Although the demarcation lines are created to represent this fractional AGN contribution, spaxels to the right may suffer from LINER contamination which can give rise to similar line ratios to that of Seyfert 2 emission. Thus, the rightmost part of the diagram is conservatively representing the broader family of LINERs, which may be powered by AGN. The region in between both demarcation lines is analogous to the composite region in traditional line diagnostic diagrams.

### v2

@@ 4 Results
The MUSE data reveals 62/126 () non-central ETGs with H emission in their spectrum, tracing warm ionized gas. These nebulae are divided into two general categories, based on their morphology and kinematics: rotating disks and filamentary networks. The histogram presented in the bottom panel of Fig. 2 shows the count and stellar mass distribution of these two morphological classifications. The full stellar mass range of all sources with H emission span 8.25 log(M⋆/M⊙) 11.17. The warm gas morphology and stellar mass of all 62 non-central ETGs are provided in Table. LABEL:tab:results.
@@ 4.1 Rotating warm gas disks
The rotating gas disks dominate the H morphological distribution in our sample. We find that 35 out of 62 H-emitters exhibit rotating disks, which we define as warm nebulae showing a clear velocity gradient indicative of ordered rotation. The sizes of the gas disks, defined as the projected length of the disk’s semi-major axis, range from roughly 0.1 - 8 kpc.
@@ 4.2 Warm gas filaments
The remaining 27 non-central ETGs with H emission have warm gas morphologies in the form of filamentary networks. A source is classified as filamentary if its gas nebulae show compact or extended, irregular structures–typically extending from the optical center of the galaxy. Additionally their gas velocity fields are generally absent of clear, ordered rotation.
The optical continuum, H flux, velocity, and velocity dispersion for all 27 filamentary sources in our sample are shown in Fig. 4. To quantify the size of the filaments, we refer to the projected distance from the optical center of the galaxy to its most radially extended filament. The size of the filaments range from 0.3 - 4 kpc.
@@ 4.3 Stellar kinematics
Using the GIST pipeline (see Sec. 3.1.1) we derive the line-of-sight velocity and velocity dispersion maps tracing the stellar population. of our ETGs with H emission show regular, ordered rotation in their stellar velocity fields. Low-mass (log(M⋆/M⊙) 9.5) sources occupy of the ETGs that are absent of ordered stellar rotation. For some of these low-mass ETGs, this may be attributed to their low stellar-surface brightness preventing the pipeline from adequately deriving the galaxy’s underlying line-of-sight velocity distribution.
The H and stellar velocity fields of ETGs hosting rotating H disks in general show kinematic and spatial overlap between their gaseous and stellar disks. This is not the case for the H filamentary sources. The asymmetric, non-rotating filamentary nebulae in many sources appear to be dissociated with their stellar population (see Sec. 5.1.1 and 5.2.4 for further discussion).
@@ 4.4 Hot gas content
In Fig. 5, we compare the diffuse X-ray luminosity in 0.5-2.0 keV (normalized by stellar mass) against the stellar mass for all ETGs in our sample with available MUSE and Chandra data. The horizontal dashed line provides the average X-ray emissivity of unresolved LMXBs from Hou et al. (2021). Galaxies sitting above this line have soft diffuse X-ray emission exceeding what is expected from unresolved LMXBs and likely contain a hot gas reservoir. We distinguish our data by those with filamentary ionized gas (red x’s), ionized gas in the form of rotating disks (blue circles), and those where H went undetected (gray diamonds). The inverted filled triangles show filamentary systems where the emissivity was computing using an upper-limit on . Inverted open triangles show filamentary systems not observed by Chandra, and their upper-limit on from eROSITA observations is gathered from eRASS1 (Predehl et al., 2021; Tubín-Arenas et al., 2024)333We homogenized the eROSITA catalog upper limits (originally reported in the 0.6-2.3 keV band, assuming an absorbed power-law spectrum with and cm-2) by computing a bandpass/model conversion factor and then re-expressing them as 0.5-2.0 keV upper limits under our fiducial thermal model., when available. All ETGs with filamentary ionized gas are likely to contain hot gas, as are a strong percentage of the ETGs containing rotating gas disks. Roughly half of the ETGs where H went undetected in emission contain hot gas. Therefore, we observe a mild connection between the hot gas content and the morphology of the warm gas, when compared at nearly the same mass range.
![[Uncaptioned image]](x5.png)
Figure 4 (continued).
![[Uncaptioned image]](x6.png)
Figure 4 (continued).
![[Uncaptioned image]](x7.png)
Figure 4 (continued).
@@ 4.5 A case study with NGC 4374
We further derived the electron pressure of each of the two phases within the same regions. To derive the pressure of the warm phase, we assume a uniform electron temperature K. The warm gas electron number density () can be computed at each spaxel using PyNeb (Luridiana et al., 2015), given that spaxel’s [SII]6717/[SII]6731 flux ratio. Note that variations in the warm gas temperature are expected, and should be considered when interpreting the derived under our isothermal assumption.
To determine the electron temperature and number density of the hot gas, we extracted the X-ray spectra from these regions (using the same local background described earlier). We then fit the spectra with an absorbed thermal plasma model, phabsapec, in Xspec version 12.12.1 (Arnaud, 1996). We derive of the hot gas using the fit’s reported normalization parameter, , via:
@@ 5.1.1 Stellar mass loss
Processes involving an ETG’s stellar population can build up its warm ISM, specifically continual mass loss from post-asymptotic giant branch stars (pAGB) and planetary nebulae (Parriott and Bregman, 2008; Bregman and Parriott, 2009; Voit and Donahue, 2011). Material that is gradually shed from stars should be mostly shock-heated and mix in with the ambient, hot atmosphere. However, numerical hydrodynamics suggest that under the right conditions pertaining to the density of the hot ambient medium, as well as a star’s velocity through that medium, can allow the shed material to dodge this shock heating thus preventing it from entering the hot phase (Parriott and Bregman, 2008). In this case, small, but non-negligible amounts of stellar mass loss will remain cool and be distributed within the ISM.
If stellar mass loss is the dominant origin of warm gas, one would expect there to be spatial and kinematic overlap between the stars and gas. This can be observed by comparing the velocity fields of both populations. To do this, we calculate the kinematic position angle (PA) of both the stars and gas using the PaFit Python package444https://pypi.org/project/pafit/ which implements the algorithm provided in the appendix of Krajnovic et al. (2006). The PA allows for a consistent, geometric interpretation of the bulk motion of the stars and gas. Specifically, it is defined as the angle sweeping from the (projected) north to the location of the maximum positive velocity, in the counterclockwise direction.
@@
Using the threshold of from Lagos et al. (2015) to classify sources as having decoupled stars and nebulae, we find that 22/27 () filamentary sources, show kinematic misalignment – unlike the majority of the rotating disks. This is consistent with our analysis in Eskenasy et al. (2024), where we found that all non-central ETGs with warm filaments had widely offset PAs between the gas and stars, albeit with a significantly smaller sample size. The frequently observed misalignment observed in ETGs with warm filaments suggests that the two populations are generally decoupled from each other in this larger sample, hinting that stellar mass loss is likely a subdominant process in forming their warm gas.
@@ 5.1.2 Cooling from the hot atmosphere
In a recent study of cool-core filaments, a close correlation has been identified between their X-ray surface brightness and H surface brightness – SBX SB (Olivares et al., 2025). Strikingly, the slope of this relation matches exactly that found for the X-ray and H surface brightnesses of stripped tails of jellyfish galaxies (Sun et al., 2022), demonstrating that the multiphase gas in both settings may share the same origin. The formation mechanism of the warm gas can occur in a broader astrophysical setting, and the presence of X-ray emitting hot gas is indispensable in this process. As shown in Fig. 5, all filamentary ETGs are consistent with hosting a detectable hot-gas component, whereas not all rotating disk sources or systems without lacking H-emission do so. In particular, only one filamentary system lies near the expected contribution from unresolved LMXBs, but when accounting for both the measurement uncertainty and uncertainty in the LMXB relation, it is not significantly inconsistent with an excess above the unresolved LMXB-only expectation. This may imply that the presence of hot gas is a necessary but not sufficient condition for the formation of filamentary warm gas. We overlay NGC 4477 in Fig. 5 (black square) as it is the smallest known system with X-ray cavities and cool-core signatures (Li et al., 2018). Although it was not observed with MUSE, it serves as a reference point demonstrating that the filamentary ETGs within our sample possess a higher hot gas content than NGC 4477555Crocker et al. (2011) found that the nucleus of NGC 4477 hosts a small-scale, rotating ring of molecular gas..
@@
from Eq. 1 in Nesvadba et al. (2007), assuming case B recombination (Osterbrock and Bochkarev, 1989). is the H luminosity (in units of erg s-1) and was extracted from the MUSE H flux maps. is the electron density in units of 100 cm-3 (determined using the procedure explained in Sec. 4.5).
@@
adopted from Su et al. (2017). Here, is the average particle weight and is the proton mass. is the bolometric X-ray luminosity, which we estimated using our derived 0.5-2.0 keV luminosities provided in Table. LABEL:tab:results and the energy conversion factor from Xspec version 12.12.1. We assume keV for all sources, with the exception of NGC 1266 and NGC 4374 where we assume keV and , respectively.
@@
Fig. 8 shows the ionized gas masses of all filamentary sources and the time required to produce such reservoirs via cooling. The crosses and inverted filled triangles show filamentary systems where was derived using Chandra detections and upper limits, respectively. Inverted open triangles show filamentary systems not observed by Chandra, and their upper-limit on is gathered from eRASS1 when available. Our derived timescales span years, all less than the Hubble time. These estimates suggest that it is indeed possible for the filamentary warm gas reservoirs in our non-central ETGs to be produced through cooling.
@@
To further investigate the cooling mechanism in non-central ETGs and its connection to BCGs and jellyfish galaxies, we performed an in depth study of NGC 4374, as it is the closest analogy of cool-core BCGs in our sample (see Sec. 4.5). We compared the X-ray to H surface brightness of the nine regions, finding a median value of , which is larger than the BCG filaments and tails of jellyfish galaxies, differing by 9.06. As described in Sec. 4.5, we derived the pressure of each gas phase and found on that, on average, the warm filaments, especially those extended to large radii, have higher electron pressure than that in the hot filaments. This result is opposite to what was found in Olivares et al. (2025) for multiphase filaments in cool-core BCGs. Given that the hot and warm gas should be roughly in pressure equilibrium with each other, Olivares et al. (2025) posited that the warm filaments have additional non-thermal pressure components. In the case of NGC 4374, additional pressure support is needed in the X-ray filaments.
@@
We investigate the role of non-thermal components such as turbulence and magnetic fields, that may be adding pressure support to NGC 4374’s hot ISM. To estimate the turbulent pressure, we refer to Ogorzalek et al. (2017) which computed the 1D turbulent velocity measurements of NGC 4374 and other massive ellipticals through line broadening and resonant scattering measurements. Interestingly, NGC 4374 was found to have one the lowest 1D velocity dispersions in their sample, with an upper limit of km s-1. The turbulent pressure () can be calculated with where is the gas mass density666 where is the mean molecular weight per electron. is the electron number density estimated in Eq. 1. is the proton mass. and , assuming isotropic turbulence. This would bring the average hot gas turbulent pressure among the regions we consider to be dyn cm-2.
@@
The magnetic pressure () can be estimated as where is the magnetic field amplitude. We refer to Finoguenov and Jones (2001) for the estimate of . In their study, they estimated the strength of the magnetic field by determining the Faraday rotation from radio observations, as well as the electron number density of the hot gas which was extracted from the hot gas’s three dimensional density profile using Chandra X-ray observations. They find that the magnetic field associated with the X-ray plasma is particularly weak at G. This corresponds to dyn cm-2. Our estimates of the non-thermal pressure support added by turbulence and magnetic fields in the hot gas still does not account for its pressure imbalance with the warm gas. We posit that an additional non-thermal pressure component. For example, cosmic rays accelerated by NGC 4374’s prominent AGN jet (Jenkins et al., 1977; Finoguenov and Jones, 2001) may be present. Although, the degree to which cosmic rays add pressure support is not well known.
@@ 5.1.3 Mergers and external accretion
Kinematic misalignment between an ETG’s warm gas and stars can be a key tracer of warm gas having an external origin such as galaxy mergers, tidal stripping from satellites, and accretion. Davis et al. (2011) found that it is common for local ETGs to exhibit this misalignment, thus demonstrating that a non-negligible fraction of warm gas in our sample of non-central ETGs is likely to be expected777Although subtle, but key differences in the merger histories -particularly timescales - arise from ranging environments, stellar masses, and whether an ETG is a fast or slow rotator.. Our comparison of the kinematic position angles in Fig. 7 shows the majority of the filamentary systems have decoupled stars and gas suggesting an external origin of their warm gas.
@@ 5.2 Powering of line emission
The traditional optical-line-diagnostic diagram proposed by Baldwin, Phillips and Terlevich (BPT; Baldwin et al., 1981) is widely used to distinguish ionization sources in galaxies. However, it is known to have a variety of systematic uncertainties which can lead to ambiguous classification of these ionizing sources. Here, we adopt a modified optical-line-diagnostic diagram from Ji and Yan (2020) to determine the ionization mechanisms in our sample. This diagram is a reprojected version of the BPT diagram, which ensures consistency of the model predictions among all three logarithmic line ratios: (), (), and (). The axes of this diagram are linear combinations of these three quantities, defined as
@@
Fig. 9 shows our systems placed in this reprojected diagram. As described in Ji and Yan (2020), spaxels that fall to the left suffer from less than 10 percent AGN contamination, while those to the right suffer from greater than 10 percent AGN contamination. Although the demarcation lines are created to represent this fractional AGN contribution, spaxels to the right may suffer from LINER contamination which can give rise to similar line ratios to that of Seyfert 2 emission. Thus, the rightmost part of the diagram is conservatively representing the broader family of LINERs, which may be powered by AGN. The region in between both demarcation lines is analogous to the composite region in traditional line-diagnostic diagrams.